OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did Connie receive? A. 53 B. 84 C. 137 D. 168

OpenStudy (anonymous):

If I am correct H=53+C 221=H+C 221=(53+C)+C Solve from there.

OpenStudy (anonymous):

So you got the answer? If you did it should be 84